A company produces a product that contains 12 ounces of materials in each unit of finished goods. During the production process, 10% of the materials evaporate. The company pays its suppliers $3 per ounce; the cost to ship the material to the company averages  $0.30 per ounce. The standard dollar amount of raw materials contained in one unit of finished goods is
Answer:
(round to nearest cents)
Answer #1

Total raw materials per unit = 12 + 12*10% = 13.2 ounces

Standard dollar amount of raw materials = 13.2*(3+0.30) = $43.56